Use elimination to solve the system of equations
4.25x - 1.28y = -9.2
x + 1.28y = 17.6

Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Use elimination to solve the system of equations
4.25x - 1.28y = -9.2
x + 1.28y = 17.6
-------
Add and solve for "x"::
5.25x = 8.4
-----
x = 1.6
-----------------
Solve for "y"::
x + 1.28y = 17.6
1.28y = 17.6-1.6
1.28y = 16
---
y = 12.5
-------------
Cheers,
Stan H.
------------

Answer by addingup(3677)

4.25x - 1.28y = -9.2
x + 1.28y = 17.6 multiply both sides of this equation times -4.25 and add to the one above:
4.25x - 1.28y = -9.2
+
-4.25x - 5.44y = -74.8
-------------------- -
. . . -6.72y = -84 divide both sides by 6.72 (remember: -/- = +)
y = 12.5 Now find x:
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
x+1.28(12.5) = 17.6
x+16 = 17.6
x = 1.6
There you have it. Check the answers by substituting in the equations, I'll do one you do the other:
4.25x-1.28y = -9.2
4.25(1.6)-1.28(12.5) = -9.2
6.8-16 = -9.2 Subtract on left to see that this is correct. You do the other equation.